# Quantum synchronization between two strongly driven YIG spheres mediated via a microwave cavity

### v3

@@ II MODEL AND EQUATION OF MOTION
Figure 2: Limit-cycle trajectories in the $\bar{q_{1}}\leftrightharpoons\bar{p_{1}}$ (red) and $\bar{q_{2}}\leftrightharpoons\bar{p_{2}}$ (blue) spaces (a, d, g), and variation of the mean values $\bar{q}_{1}$ (red) and $\bar{q}_{2}$ (blue) (b, e, h), along with mean values $\bar{p}_{1}$ (red) and $\bar{p}_{2}$ (blue) (c, f, i). The chosen parameters are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=g_{2}=0.1$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $g_{1}=g_{2}=0.1$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=0.2$, $g_{2}=0.1$. All other parameters are identical across cases: $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\bar{q}_{1},\bar{p}_{1})=(1,0)=(\bar{q}_{2},\bar{p}_{2})$ [i.e., $\alpha_{i}=1/\sqrt{2}$ well satisfying the mean-field approximation]. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 3: Limit-cycle trajectories in the $\bar{q_{1}}\leftrightharpoons\bar{p_{1}}$ (red) and $\bar{q_{2}}\leftrightharpoons\bar{p_{2}}$ (blue) spaces (a, d, g); variation of the mean values $\bar{q}_{1}$ (red), $\bar{q}_{2}$ (blue) (b, e, h); and $\bar{p}_{1}$ (red), $\bar{p}_{2}$ (blue) (c, f, i). The parameters for each case are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$$\Delta_{1}=\Delta_{2}=0.001$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Delta_{1}=\Delta_{2}=0.001$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $\Delta_{1}=0.001$, $\Delta_{2}=0.0025$. Other parameters (common to all cases): $\Omega_{c}=1$, $\Delta_{c}=-0.2$, $g_{1}=g_{2}=0.5$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\bar{q}_{1},\bar{p}_{1})=(1,0)$ and $(\bar{q}_{2},\bar{p}_{2})=(2,0)$ [i.e., $\alpha_{1}=1/\sqrt{2}$ and $\alpha_{2}=\sqrt{2}$ respectively]. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 4: (a) Variation of $S_{q}^{\phi}$, with respect to time $t$ for $\bar{n}_{m}=0$ (blue solid) and $\bar{n}_{m}=0.1$ (red dashed), (b) variation of $\bar{S}_{q}^{\phi}$, with respect to the mean phonon number $\bar{n}_{m}$ of the environment. The other parameters are $g_{1}=g_{2}=0.1$, $K_{1}=K_{2}=10^{-10}$, $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$, and $\phi=0.1320$ radian. We have chosen an initial condition $(\bar{q}_{1},\bar{p}_{1})=(1,1)=(\bar{q}_{2},\bar{p}_{2})$ [i.e., $\alpha_{i}=e^{\iota\pi/4}$]
@@ II.1 Solution in mean-field approximation
In the limit of negligible higher-order fluctuations, we obtain the linearized equations for the quantum fluctuations, as well, given by
where $\delta m_{i}=\frac{\delta q_{i}+\iota\,\delta p_{i}}{\sqrt{2}},\quad i\in 1,2$  and $\delta a=\frac{\delta x+\iota\,\delta y}{\sqrt{2}}$ . The expressions of the parameters $P_{i}$, $Q_{i}$, $R_{i}$, $S_{i}$, $U_{i}$, $W_{i}$, and $T$ are listed in Appendix A.
To represent the magnon dynamics in phase space, we need to introduce the relevant quadratures as follows: $q_{i}=\frac{1}{\sqrt{2}}\left(m_{i}^{\dagger}+m_{i}\right)$, $p_{i}=\frac{\iota}{\sqrt{2}}\left(m_{i}^{\dagger}-m_{i}\right)$, $x=\frac{1}{\sqrt{2}}(a^{\dagger}+a)$ and $y=\frac{\iota}{\sqrt{2}}(a^{\dagger}-a)$, where $q_{i}$ and $p_{i}$ are the position and linear momentum quadratures of the $i$th magnon mode.
To ascertain the degree of quantum synchronization between oscillators, we adopt a figure of merit originally proposed by Mari et al. PhysRevLett.111.103605 . In the case when the quantum fluctuations can be ignored, the classical synchronization prevails and may be quantified as
@@
are the differences between the position quadratures $q_{1,2}$ of the two oscillators and between their linear momentum quadratures $p_{1,2}$, respectively. In the limit, when $q_{-},p_{-}\rightarrow 0$, the $S_{c}\rightarrow\infty$ denotes complete classical synchronization. In phase space, they would represent a limit cycle, with a constant phase difference between them. Note that, in this analysis, the quadratures are defined in terms of their mean values $\bar{q}_{i}$ and $\bar{p}_{i}$, respectively.
In the quantum regime, when quantum fluctuations cannot be ignored, the above variables need to be redefined with respect to their mean values as follows:
$q_{-}(t)$ $\rightarrow q_{-}(t)-\bar{q}_{-}(t)=\delta q_{-}(t)\;,$
$p_{-}(t)$ $\rightarrow p_{-}(t)-\bar{p}_{-}(t)=\delta p_{-}(t)\;.$ (16)
The quantum synchronization measure can then be expressed only in terms of these fluctuations $\delta q_{-}$ and $\delta p_{-}$ as
@@
and $\phi_{i}=\tan^{-1}\left[\frac{\bar{p}_{i}(t)}{\bar{q}_{i}(t)}\right]$. Therefore, the generalized form of the Eq. (17) - the so-called “quantum $\phi$ synchronization” qiao2020quantum ; sun2024quantum - can be defined as
@@
Note that the above quantifier of synchronization is a marker of classical phase locked complete quantum synchronization and not ‘quantum phase synchronization’. Within this definition, $S_{q}^{\phi}(t)$ is invariant under arbitrary choices of the constant phase offset $\phi$, and therefore depends only on the relative phase locking between the two limit cycles rather than on the absolute phase reference. To calculate $S_{q}^{\phi}(t)$, we need to solve for the quadrature fluctuations of the oscillators. Therefore, it becomes more convenient to obtain the corresponding equations from Eqs. (II.1), by replacing the relevant operators and input noise operators with their quadratures: $\delta q_{i}=\frac{1}{\sqrt{2}}\left(\delta m^{\dagger}_{i}+\delta m_{i}\right)$, $\delta p_{i}=\frac{\iota}{\sqrt{2}}\left(\delta m^{\dagger}_{i}-\delta m_{i}\right)$, $q_{in}=\frac{1}{\sqrt{2}}\left(m_{in}^{\dagger}+m_{in}\right)$, $p_{in}=\frac{\iota}{\sqrt{2}}\left(m_{in}^{\dagger}-m_{in}\right)$ , $\delta x=\frac{1}{\sqrt{2}}(\delta a^{\dagger}+\delta a)$, $\delta y=\frac{i}{\sqrt{2}}(\delta a^{\dagger}-\delta a)$, $x_{in}=\frac{1}{\sqrt{2}}(a_{in}^{\dagger}+a_{in})$, and $y_{in}=\frac{i}{\sqrt{2}}(a_{in}^{\dagger}-a_{in})$.
The equations for quadrature fluctuations then take a consolidated form as given by
@@
$\mathbf{M}(t)=\begin{pmatrix}P_{1}&Q_{1}&0&0&R_{1}&S_{1}\\ Q_{1}^{*}&P_{1}^{*}&0&0&S_{1}^{*}&R_{1}^{*}\\ 0&0&P_{2}&Q_{2}&R_{2}&S_{2}\\ 0&0&Q_{2}^{*}&P_{2}^{*}&S_{2}^{*}&R_{2}^{*}\\ U_{1}&W_{1}&U_{2}&W_{2}&T&0\\ W_{1}^{*}&U_{1}^{*}&W_{2}^{*}&U_{2}^{*}&0&T^{*}\end{pmatrix}$, (23)
@@ III Results
In this section, we will discuss by numerically solving the Eqs. (11), (12), and (II.1), how the quantum synchronization between two magnon modes is developed. We consider a strong driving regime, when the Rabi frequencies $\Omega_{i}$ and $\Omega_{c}$ of the driving fields are much larger than the decay rates $\gamma_{i}$ of the $i$th magnon mode and $\gamma_{c}$ of the cavity mode, respectively. This regime would otherwise render an independent dynamics of the magnon mode in the absence of the cavity mode, exhibiting an underdamped oscillation that could sustain up to a time scale $t\sim 1/\gamma_{i}$. In the present case, the cavity mode creates an indirect coupling between the magnon modes, leading to classical synchronization between them that persists up to a timescale much larger than $1/\gamma_{i}$. We show these results in Figs. 2.
We first choose $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}\Omega_{1}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\bar{q}_{1}$, $\bar{q}_{2}$ and $\bar{p}_{1}$, $\bar{p}_{2}$ become indistinguishable at times as long as $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of two YIG spheres have been previously explored in a cavity-magnomechanical setup, in which the phase synchronization has been achieved between the vibrational degrees of freedom, instead of magnon modes PhysRevResearch.5.043197 . When the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, the system still maintains the phase synchronization, with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f) between the trajectories of the two magnon modes, although the quadratures still oscillate at the same frequency. Such a behavior refers to a classical phase synchronization.
@@
One of the important features of classical synchronization is its robustness against a large change in initial conditions. We verify the existence of limit cycles by varying the initial conditions (Fig. 3a, 3b, and 3c). We choose an initial condition $\bar{q}_{1}(0),\bar{p}_{1}(0)]=[1,0]$ and $\bar{q}_{2}(0),\bar{p}_{2}(0)]=[2,0]$. For $\Delta\Omega=10^{-5}\Omega_{1}$, both magnon modes rapidly converge to a common limit cycle in phase space. Similar feature persists also when the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, though with a finite phase difference of $\phi=1.2276$ radian (Figs. 3d, 3e, and 3f). In this regime, the system exhibits chaotic transients before settling into a shared limit cycle, providing a clear signature of synchronization. Furthermore, when the detunings of the driving fields are chosen differently, e.g., $\Delta_{1}=0.001$ and $\Delta_{2}=0.0025$, the system continues to display phase synchronization with an increased phase difference of $\phi=1.5337$ radian (Figs. 3g, 3h, and 3i).
@@
In contrast Figure 4a, for a small thermal population $\bar{n}_{m}=0.1$, $S_{q}^{\phi}$ still grows monotonically but saturates at a reduced value $S_{q}^{\phi}\simeq 0.83$, reflecting partial degradation of synchronization due to thermal noise. Nevertheless, in both cases the system reaches a steady synchronized regime, demonstrating that cavity-mediated magnon synchronization remains robust against weak thermal decoherence.

### v4

# arXiv:2507.17042v4[quant-ph] 01 Mar 2026

@@ II MODEL AND EQUATION OF MOTION
Figure 2: Limit-cycle trajectories in the $\langle{q_{1}\rangle}\leftrightharpoons\langle{p_{1}\rangle}$ (red) and $\langle{q_{2}\rangle}\leftrightharpoons\langle{p_{2}\rangle}$ (blue) spaces (a, d, g), and variation of the mean values $\langle{q}_{1}\rangle$ (red) and $\langle{q}_{2}\rangle$ (blue) (b, e, h), along with mean values $\langle{p}_{1}\rangle$ (red) and $\langle{p}_{2}\rangle$ (blue) (c, f, i). The chosen parameters are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=g_{2}=0.1$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $g_{1}=g_{2}=0.1$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $g_{1}=0.2$, $g_{2}=0.1$. All other parameters are identical across cases: $\Omega_{c}=1$, $\Delta_{1}=\Delta_{2}=0.001$, $\Delta_{c}=-0.2$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\langle{q}_{1}\rangle,\langle{p}_{1}\rangle)=(1,0)=(\langle{q}_{2}\rangle,\langle{p}_{2}\rangle)$ [i.e., $\alpha_{i}=1/\sqrt{2}$ well satisfying the mean-field approximation]. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 3: Limit-cycle trajectories in the $\langle{q_{1}\rangle}\leftrightharpoons\langle{p_{1}\rangle}$ (red) and $\langle{q_{2}\rangle}\leftrightharpoons\langle{p_{2}\rangle}$ (blue) spaces (a, d, g); variation of the mean values $\langle{q}_{1}\rangle$ (red), $\langle{q}_{2}\rangle$ (blue) (b, e, h); and $\bar{p}_{1}$ (red), $\langle{p}_{2}\rangle$ (blue) (c, f, i). The parameters for each case are: (a–c) $\Omega_{1}=1$, $\Omega_{2}=1.00001$$\Delta_{1}=\Delta_{2}=0.001$. (d–f) $\Omega_{1}=1$, $\Omega_{2}=1.1$, $\Delta_{1}=\Delta_{2}=0.001$. (g–i) $\Omega_{1}=1$, $\Omega_{2}=1.00001$, $\Delta_{1}=0.001$, $\Delta_{2}=0.0025$. Other parameters (common to all cases): $\Omega_{c}=1$, $\Delta_{c}=-0.2$, $g_{1}=g_{2}=0.5$, $K_{1}=K_{2}=10^{-10}$, and $\gamma_{1}=\gamma_{2}=\gamma_{c}=0.1$. All parameters are normalized with respect to $\Omega_{1}$ with the initial conditions $(\langle{q}_{1}\rangle,\langle{p}_{1}\rangle)=(1,0)$ and $(\langle{q}_{2}\rangle,\langle{p}_{2}\rangle)=(2,0)$ [i.e., $\alpha_{1}=1/\sqrt{2}$ and $\alpha_{2}=\sqrt{2}$ respectively]. The limit cycle trajectories are shown till $\Omega_{1}t=10^{5}$.
Figure 4: Time evolution of the fluctuation-to-mean ratios $\langle\delta m_{i}^{\dagger}\delta m_{i}\rangle/|\alpha_{i}|^{2}$ ($i=1,2$) and $\langle\delta a^{\dagger}\delta a\rangle/|\beta|^{2}$, for the parameters as in Figs. 2.
@@ II.1 Solution in mean-field approximation
To represent the magnon dynamics in phase space, we introduce the quadrature operators $q_{i}=\frac{1}{\sqrt{2}}(m_{i}^{\dagger}+m_{i})$ and $p_{i}=\frac{\iota}{\sqrt{2}}(m_{i}^{\dagger}-m_{i})$ for the $i$th magnon mode, and for the cavity mode $x=\frac{1}{\sqrt{2}}(a^{\dagger}+a)$ and $y=\frac{\iota}{\sqrt{2}}(a^{\dagger}-a)$. These operators define the canonical phase-space variables of the corresponding cavity modes. The expectation values of the quadratures, $\langle q_{i}\rangle$ and $\langle p_{i}\rangle$, determine the real and imaginary parts of $\alpha_{i}$, respectively, such that $\alpha_{i}=\left(\langle q_{i}\rangle+\iota\langle p_{i}\rangle\right)/\sqrt{2}$. Analogously, the cavity amplitude is expressed in terms of its phase-space coordinates as $\beta=\left(\langle x\rangle+\iota\langle y\rangle\right)/\sqrt{2}$. Thus, the complex amplitudes $\alpha_{i}$ and $\beta$ are uniquely specified by the corresponding quadrature expectation values, providing a direct geometrical interpretation of the mean-field dynamics in phase space.
In the case when the quantum fluctuations can be ignored, the classical synchronization prevails and may be quantified as
@@
are the differences between the position quadratures $q_{1,2}$ of the two oscillators and between their linear momentum quadratures $p_{1,2}$, respectively. In the limit, when $q_{-},p_{-}\rightarrow 0$, the $S_{c}\rightarrow\infty$ denotes complete classical synchronization. In phase space, they would represent a limit cycle, with a constant phase difference between them.
II.2 Quantum regime
To describe the dynamics beyond the semiclassical approximation, we retain the quantum fluctuations around the steady-state mean values and neglect higher-order nonlinear terms in the fluctuation operators. Therefore, the equations of motion for the fluctuation operators take the form
The expressions of the parameters $P_{i}$, $Q_{i}$, $R_{i}$, $S_{i}$, $U_{i}$, $W_{i}$, and $T$ are listed in Appendix A.
To ascertain the degree of quantum synchronization between oscillators, we adopt a figure of merit originally proposed by Mari et al. PhysRevLett.111.103605 . In this regime, when quantum fluctuations cannot be ignored, the quadrature variables need to be redefined with respect to their expectation values as follows:
$q_{-}(t)$ $\rightarrow q_{-}(t)-\langle q_{-}(t)\rangle=\delta q_{-}(t)\;,$
$p_{-}(t)$ $\rightarrow p_{-}(t)-\langle p_{-}(t)\rangle=\delta p_{-}(t)\;.$ (16)
The quantum synchronization measure can then be expressed only in terms of these fluctuations $\delta q_{-}=\delta q_{1}-\delta q_{2}$ and $\delta p_{-}=\delta p_{1}-\delta p_{2}$ as
@@ II.2 Quantum regime
Here, we define the dimensionless quadrature fluctuations as $\delta q_{i}=\frac{1}{\sqrt{2}}\left(\delta m_{i}^{\dagger}+\delta m_{i}\right)$ and $\delta p_{i}=\frac{\iota}{\sqrt{2}}\left(\delta m_{i}^{\dagger}-\delta m_{i}\right)$.
@@
and $\phi_{i}=\tan^{-1}\left[\frac{\langle{p}_{i}\rangle(t)}{\langle{q}_{i}\rangle(t)}\right]$. Therefore, the generalized form of the Eq. (17) - the so-called “quantum $\phi$ synchronization” qiao2020quantum ; sun2024quantum - can be defined as
@@
Note that the above quantifier of synchronization is a marker of classical phase-locked complete quantum synchronization, and not ‘quantum phase synchronization’. Within this definition, $S_{q}^{\phi}(t)$ is invariant under arbitrary choices of the constant phase offset $\phi$, and therefore depends only on the relative phase locking between the two limit cycles rather than on the absolute phase reference.
To calculate $S_{q}^{\phi}(t)$, we need to solve for the quadrature fluctuations of the oscillators. Therefore, it becomes more convenient to obtain the corresponding equations from Eqs. (II.2), by replacing the relevant fluctuation operators and input noise operators with their quadratures, namely, $\delta x=\frac{1}{\sqrt{2}}(\delta a^{\dagger}+\delta a)$, $\delta y=\frac{\iota}{\sqrt{2}}(\delta a^{\dagger}-\delta a)$, $q_{\rm in}=\frac{1}{\sqrt{2}}\left(m_{\rm in}^{\dagger}+m_{\rm in}\right)$, $p_{\rm in}=\frac{\iota}{\sqrt{2}}\left(m_{\rm in}^{\dagger}-m_{\rm in}\right)$, $x_{\rm in}=\frac{1}{\sqrt{2}}(a_{\rm in}^{\dagger}+a_{\rm in})$, and $y_{\rm in}=\frac{\iota}{\sqrt{2}}(a_{\rm in}^{\dagger}-a_{\rm in})$. The equations for quadrature fluctuations then take a consolidated form as given by
@@
[…]
The vector $N(t)$ containing the noise terms is given below:
@@
$M(t)=\begin{pmatrix}P_{1}&Q_{1}&0&0&R_{1}&S_{1}\\ Q_{1}^{*}&P_{1}^{*}&0&0&S_{1}^{*}&R_{1}^{*}\\ 0&0&P_{2}&Q_{2}&R_{2}&S_{2}\\ 0&0&Q_{2}^{*}&P_{2}^{*}&S_{2}^{*}&R_{2}^{*}\\ U_{1}&W_{1}&U_{2}&W_{2}&T&0\\ W_{1}^{*}&U_{1}^{*}&W_{2}^{*}&U_{2}^{*}&0&T^{*}\end{pmatrix}$, (23)
@@ III Results
In this section, we will discuss by numerically solving the Eqs. (11), (12), and (22), how the quantum synchronization between two magnon modes is developed. We consider a strong driving regime, when the Rabi frequencies $\Omega_{i}$ and $\Omega_{c}$ of the driving fields are much larger than the decay rates $\gamma_{i}$ of the $i$th magnon mode and $\gamma_{c}$ of the cavity mode, respectively. This regime would otherwise render an independent dynamics of the magnon mode in the absence of the cavity mode, exhibiting an underdamped oscillation that could sustain up to a time scale $t\sim 1/\gamma_{i}$. In the present case, the cavity mode induces an indirect coupling between the magnon modes, leading to classical synchronization between them that persists on a timescale much longer than $1/\gamma_{i}$. We show these results in Figs. 2.
We first choose $\Delta\Omega=\Omega_{2}-\Omega_{1}=10^{-5}\Omega_{1}$, for which both the magnon modes rapidly converge to a shared limit cycle in phase space (Figs. 2a2b, and 2c). This behavior is a clear signature of classical synchronization, where the mean quadratures $\langle{q}_{1}\rangle$, $\langle{q}_{2}\rangle$ and $\langle{p}_{1}\rangle$, $\langle{p}_{2}\rangle$ become indistinguishable at times as long as $t\approx 10^{5}/\Omega_{i}\gg 1/\gamma_{i}$. Such phase-locked dynamics of two YIG spheres have been previously explored in a cavity-magnomechanical setup, in which the phase synchronization has been achieved between the vibrational degrees of freedom, instead of magnon modes PhysRevResearch.5.043197 . When the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, the system still maintains the phase synchronization, with a nonzero phase difference $\phi=0.1320$ radian (Figs. 2d2e, and 2f) between the trajectories of the two magnon modes. However, the quadratures still oscillate at the same frequency. Such a behavior refers to a classical phase synchronization.
@@
One of the important features of classical synchronization is its robustness against a large change in initial conditions. We verify the existence of limit cycles by varying the initial conditions (Fig. 3a, 3b, and 3c). We choose an initial condition $\langle{q}_{1}\rangle(0),\langle{p}_{1}\rangle(0)]=[1,0]$ and $\langle{q}_{2}\rangle(0),\langle{p}_{2}\rangle(0)]=[2,0]$. For $\Delta\Omega=10^{-5}\Omega_{1}$, both magnon modes rapidly converge to a common limit cycle in phase space. Similar feature persists also when the Rabi frequency $\Omega_{2}$ is increased so that $\Delta\Omega=0.1\Omega_{1}$, though with a finite phase difference of $\phi=1.2276$ radian (Figs. 3d, 3e, and 3f). In this regime, the system exhibits chaotic transients before settling into a shared limit cycle, providing a clear signature of synchronization. Furthermore, when the detunings of the driving fields are chosen differently, e.g., $\Delta_{1}=0.001$ and $\Delta_{2}=0.0025$, the system continues to display phase synchronization with an increased phase difference of $\phi=1.5337$ radian (Figs. 3g, 3h, and 3i).
We have further examined the validity of the semiclassical (mean-field) approximation by quantitatively comparing the coherent mean occupations with the corresponding quantum fluctuation contributions. Specifically, we have evaluated the ratios $R_{m_{i}}=\langle\delta m_{i}^{\dagger}\delta m_{i}\rangle/|\alpha_{i}|^{2},R_{a}=\langle\delta a^{\dagger}\delta a\rangle/|\beta|^{2}$ and display their time-dependence in Fig. 4 for the same parameters as used in Fig. 2. Note that the second-order moments of the fluctuation operators can be written as linear combinations of the elements of the covariance matrix $C$, as $\langle\delta m_{1}^{\dagger}\delta m_{1}\rangle=\frac{1}{2}\left(C_{11}+C_{22}-1\right)$, $\langle\delta m_{2}^{\dagger}\delta m_{2}\rangle=\frac{1}{2}\left(C_{33}+C_{44}-1\right),$ and $\langle\delta a^{\dagger}\delta a\rangle=\frac{1}{2}\left(C_{55}+C_{66}-1\right)$. Clearly, at the long time-scale, when the quantum synchronization sets in (as we show next), these ratios remain much smaller than unity, e.g., $R_{m_{1}}=R_{m_{2}}=2.4769\times 10^{-7}$, and $R_{a}=7.46\times 10^{-3}$, thereby justifying the mean-field approximation.
@@
In contrast, for a small thermal population $\bar{n}_{m}=0.1$, $S_{q}^{\phi}$ still grows monotonically but saturates at a reduced value $S_{q}^{\phi}\simeq 0.83$, reflecting partial degradation of synchronization due to thermal noise. Nevertheless, in both cases the system reaches a steady synchronized regime, demonstrating that cavity-mediated magnon synchronization remains robust against weak thermal decoherence.